# CoPL: Collaborative Preference Learning for Personalizing LLMs

### v1

@@ 5.2 Results
Table 1 presents accuracy for both seen and unseen users. CoPL consistently outperforms other baselines, except for Oracle, in both seen user and unseen user experiments. Notably, CoPL is comparable with Oracle in UF-P-4-ALL/AVG. In unseen user experiments, CoPL achieves accuracy comparable to the seen user setting, indicating the effectiveness of our unseen user adaptation.
Fig. 1 visualizes the embedding space of seen users in UF-P-4-AVG, which is the most challenging environment in these experiments, and demonstrates that GNN-based representation learning can capture preference similarity between users even when each user provides few annotations.
@@ 5.3 Analysis
In Table 1, all models appear capable of representing diverse preferences, surprisingly including the uniform models in UF-P-2-ALL/AVG. To investigate further, we divide the test pairs of UF-P-2 into common and controversial categories, where common pairs have identical annotations from both preference groups, and controversial pairs differ. Focusing on the seen user results in UF-P-2-ALL with gemma-2b-it from Table 1, we break down the accuracy in Table 2. The results indicate that baselines, except Oracle, struggle with controversial pairs, suggesting a tendency to capture only the common preference across all users. By contrast, our method achieves comparable performance to Oracle on controversial pairs while preserving high accuracy on common pairs.
@@
Fig. 2 shows accuracy as the number of provided annotations increases in UF-P-2-AVG and UF-P-4-AVG. We observe that additional annotations lead to more accurate preference predictions for unseen users in general. However, in practice, even eight annotations are sufficient, enabling accurate inference of each user’s preference. We also compare two-hop and four-hop adaptations, but there is no significant difference.
Figure 3: Expert allocation at layers 2 and 3 in UF-P-4-ALL with gemma-2b-it. Colors indicate preference groups. Users with similar preference groups are mapped to the same expert.
UF-P-2-ALL UF-P-4-ALL
@@
Table 3: Ablation study of CoPL in UF-P-2/4-ALL with gemma-2b-it. w/o GNN embedding replaces user embeddings from GNN with learnable user embeddings. w/o MoLE removes the MoLE and projects user embeddings into the token space. The symbol $n$ denotes the LoRA rank. All experiments run on three seeds.
Ablation study of CoPL.
Table 3 presents an ablation study of CoPL, focusing on GNN-derived user embeddings and the MoLE architecture. When GNN embeddings are removed, user representations become learnable parameters. Without MoLE, user embeddings are projected into the token space and passed as an additional token to the reward model. The results indicate that components of CoPL are effective. Specifically, GNN-based embeddings are a crucial component of CoPL, and the MoLE architecture further enhances accuracy. Notably, CoPL uses fewer activated parameters than w/o MoLE $(n=64)$.
Fig. 3 depicts expert allocation across layers two and three, where the user-conditioned gating mechanism partitions users differently at each layer. We can observe that users with the same preferences tend to be routed to the same expert.
UF-P-4-ALL UF-P-4-AVG
@@
Table 4: Accuracy of unseen-user adaptation in UF-P-4-ALL/AVG with gemma-2b-it. Naive Avg. computes the unseen user’s embedding as the unweighted average of 2-hop neighbors, while CoPL applies a weighted average strategy. User Opt. represents an optimization-based approach that learns a parameterized user embedding by maximizing the likelihood of the given annotations. All experiments run on three seeds.
@@
We conduct an ablation study to evaluate the effectiveness of the unseen user adaptation strategy, comparing it to two baselines, Naive Avg and User Opt. Naive Avg assigns each unseen user embedding as the unweighted average of 2-hop seen user embeddings. User Opt replaces $e_{u}^{(L)}$ with a parameterized embedding learned by minimizing Equation (5) on the provided annotations. Table 4 reports results in UF-P-4-ALL/AVG with gemma-2b-it, showing that CoPL outperforms both alternatives while achieving better computational efficiency than the optimization-based User Opt.
Fig. 4 illustrates that naive averaging places unseen users away from identical preference group users, whereas our method clusters them more closely with users who share the same preferences.
@@
Table 5: Test accuracy of the GNN. We evaluate the model using the same users from training but with annotation pairs that are not reflected in the graph. All experiments run on three seeds.
UF-P-2-ALL UF-P-4-ALL
@@
Pseudo label $62.77_{\pm 0.70}$ $62.26_{\pm 0.27}$
Oracle $64.53_{\pm 0.14}$ $61.52_{\pm 0.13}$
User-specific $58.09_{\pm 1.73}$ $55.30_{\pm 3.30}$
Table 6: Accuracy of reward model trained by using a pre-trained GNN in UF-P-2/4-ALL with gemma-2b-it. The pseudo-label variant trains a reward model on all seen user–response pairs, with annotations provided by GNN-predicted labels. The user-specific variant refers to a BTL model trained with pseudo labels for each user. Only 10 users per group are sampled due to computational constraints. All experiments run on three seeds.
Training reward models with GNN.
Table 5 reports GNN accuracy on seen users and responses for test pairs excluded from the training dataset. The results demonstrate that GNN can accurately predict labels for unannotated pairs with sparse annotations.
Table 6 examines the impact of training with GNN-based pseudo labels, allowing the model to leverage additional preference data. Although the pseudo-labeled pairs increase the dataset size, performance is slightly worse than using only user-provided annotations, suggesting that noise degrades model accuracy.
To investigate the effect of noise further, a user-specific reward model is trained on pseudo labels for a random sample of 10 users per group. The results are considerably worse than the Oracle, indicating that noisy labels introduce training instability. This observation aligns with Wang et al. (2024), which notes that noisy preference labels can lead to training instability and performance degradation.
@@ 6 Conclusion
This work demonstrates how GCF-based user embeddings enable personalization in sparse settings, but we do not explore other GNN architectures that could further reduce sample complexity. Additionally, although CoPL employs a gating mechanism for user-specific expert allocation, we did not apply load-balancing loss, which induces more even activation among experts. As a result, some experts remain inactive in Fig. 3. Future work may investigate different GNN designs and incorporate load-balancing techniques to fully leverage the potential of GNN and MoLE, respectively.
The oracle model may appear underwhelming, likely because our smaller backbone LLM struggles to capture subtle stylistic differences between responses. Larger-scale models (over 30B parameters) could better handle these nuances; however, constraints in our current setup prevent such experiments, and we defer them to future work.
@@ Appendix B Method Baselines
Appendix C Experimental Details
@@
C.1 Ultrafeedback-P
@@
UF-P-2.
This version employs only two attributes and removes pairs that both user groups label identically, focusing on controversial cases where preferences differ.
UF-P-4.
All four attributes are retained as preference dimensions, which allows for partial agreement among groups and hence increases complexity. Although Poddar et al. (2024) also excludes pairs fully agreed upon by all users, the remaining set is larger and exhibits more variety than UF-P-2.
@@
C.2 Hyper-parameters
@@
The model consists of four message-passing layers, each with user and response embeddings of dimension 512. We use Leaky ReLU as non-linear activation function to update user and response embeddings. Training proceeds for 300 epochs using the AdamW optimizer Loshchilov (2017) with a learning rate of $1\times 10^{-4}$ and a cosine scheduler with warmup ratio $0.1$. The batch size is 1024, and all experiments are conducted on an RTX 4090 GPU.
@@
Baseline models use LoRA with rank 64. They also trained with an AdamW optimizer and a cosine scheduler with a warmup ratio $0.03$. We search the learning rate from $[1\times 10^{-4},5\times 10^{-5},1\times 10^{-5},5\times 10^{-6}]$.

### v2

@@ 5.2 Results
Table˜2 presents accuracy for both seen and unseen users. CoPL consistently outperforms other baselines, except for G-Oracle, in both seen user and unseen user experiments. Notably, CoPL surpasses the performance of G-Oracle on TL;DR and UF-P-4, demonstrating the advantage of multi-task learning. In the PersonalLLM, CoPL remains robust across the ALL and AVG, whereas VPL suffers from performance degradation in a more realistic AVG setting. These findings are consistent with Ju et al. (2024), which theoretically shows that message-passing can help users with limited interactions in collaborative filtering. In unseen user experiments, CoPL achieves accuracy comparable to the seen user setting, indicating the effectiveness of our unseen user adaptation.
Fig.˜1 illustrates the learned user embeddings for UF-P-4 (AVG), selected as the most challenging environment among those with distinct groups. The figure shows that GNN-based representation learning successfully captures preference similarities, despite the limited annotations per user.
@@ 5.3 Analysis
In Table˜2, all models appear capable of representing diverse preferences, surprisingly including the uniform models in UF-P-2 (ALL/AVG). To investigate further, we divide the test pairs of UF-P-2 into common and controversial categories, where common pairs have identical annotations from both preference groups, and controversial pairs differ. Focusing on the seen user results in UF-P-2 (ALL) with gemma-2b-it from Table˜2, we break down the accuracy in Table˜3. The results indicate that baselines, except G-Oracle, struggle with controversial pairs, suggesting a tendency to capture only the common preference across all users. By contrast, our method achieves comparable performance to G-Oracle on controversial pairs while preserving high accuracy on common pairs.
Performance under imbalanced group distributions.
We vary the group proportion from 1:9 to 9:1 on the TL;DR (AVG) and UF-P-2 (AVG) datasets. As shown in Fig.˜4, CoPL consistently captures both majority and minority preferences, maintaining stable accuracy for the short- and long-summary groups on TL;DR. On UF-P-2, CoPL still reflects diverse preferences, but the gap relative to the balanced 5:5 setting widens as the ratio becomes more skewed. Majority accuracy rises while minority accuracy falls, showing majority bias under imbalance. The lower absolute accuracy for the honesty group reflects the inherent difficulty of that preference, which remains consistent with the G-oracle results. The difference between TL;DR and UF-P-2 is also explained by UF-P-2 containing common pairs on which both groups agree, which reduces the distinct signal from the minority.
(a) TL;DR (AVG)
(b) UF-P-2 (AVG)
Figure 4: Group-wise accuracy of reward models with Gemma-2b-it in TL;DR (AVG) and UF-P-2 (AVG), varying the ratio of group size (A:B) from 1:9 to 9:1 with the total number of users fixed at 10,000.
Figs.˜A4 and A5 show the learned user embeddings for TL;DR and UF-P-2. CoPL preserves well-separated clusters aligned with group identities even under extreme imbalance. Thus, while minority group accuracy may drop, the representation space remains robust to group structure. To mitigate majority bias, we can apply loss reweighting, such as focal loss (Lin et al., 2017; Subramanian et al., 2021), to emphasize underrepresented groups in the reward model training.
In the four-group setting, additional UF-P-4 (AVG) results exhibit the same trend, as reported in Appendix˜E.
@@
Fig.˜5 shows accuracy as the number of provided annotations increases in UF-P-2 (AVG) and UF-P-4 (AVG). We observe that additional annotations lead to more accurate preference predictions for unseen users in general. However, in practice, even eight annotations are sufficient, enabling accurate inference of each user’s preference. We also compare two-hop and four-hop adaptations, but there is no significant difference.
(a) UF-P-2 (AVG)
(b) UF-P-4 (AVG)
Figure 5: Accuracy of unseen user adaptation as the number of provided annotation sets increases, evaluated on UF-P-2/4 (AVG) with gemma-2b-it. 2-hop and 4-hop indicates 2-hop and 4-hop adaptation, respectively.
Figure 6: Expert allocation at layers 2 and 3 in UF-P-4 (ALL) with gemma-2b-it. Colors indicate preference groups. Users with similar preference groups are mapped to the same expert.
UF-P-2 (ALL) UF-P-4 (ALL)
@@
Table 4: Ablation study of CoPL in UF-P-2/4 (ALL) with gemma-2b-it. w/o GNN embedding replaces user embeddings from GNN with learnable user embeddings. w/o MoLE removes the MoLE and projects user embeddings into the token space. The symbol $n$ denotes the LoRA rank.
UF-P-4 (ALL) UF-P-4 (AVG)
@@
Table 5: Accuracy of unseen-user adaptation in UF-P-4 (ALL/AVG) with gemma-2b-it. Naive Avg. computes the unseen user’s embedding as the unweighted average of 2-hop neighbors, while CoPL applies a weighted average. User Opt. represents an optimization-based approach that learns a parameterized user embedding by maximizing the likelihood of the given annotations.
Ablation study of CoPL.
Table˜4 presents an ablation study of CoPL, focusing on GNN-derived user embeddings and the MoLE architecture. When GNN embeddings are removed, user representations become learnable parameters. Without MoLE, user embeddings are projected into the token space and passed as an additional token to the reward model. The results indicate that components of CoPL are effective. Specifically, GNN-based embeddings are a crucial component of CoPL, and the MoLE architecture further enhances accuracy. Notably, CoPL uses fewer activated parameters than w/o MoLE $(n=64)$.
Fig.˜6 depicts expert allocation across layers two and three, where the user-conditioned gating mechanism partitions users differently at each layer. We can observe that users with the same preferences tend to be routed to the same expert.
We provide the ablation study of the number of experts in Appendix˜E.
@@
We conduct an ablation study to evaluate the effectiveness of the unseen user adaptation strategy, comparing it to two baselines, Naive Avg and User Opt. Naive Avg assigns each unseen user embedding as the unweighted average of 2-hop seen user embeddings. User Opt replaces $e_{u}^{(L)}$ with a parameterized embedding learned by minimizing Equation (5) on the provided annotations. Table˜5 reports results in UF-P-4-ALL/AVG with gemma-2b-it, showing that CoPL outperforms both alternatives while achieving better computational efficiency than the optimization-based User Opt.
Fig.˜A2 illustrates that naive averaging places unseen users away from identical preference group users, whereas our method clusters them more closely with users who share the same preferences.
(a) UF-P-2 (ALL)
(b) UF-P-4 (ALL)
Figure 7: Accuracy of reward models on UF-P-2 and UF-P-4 (ALL) with gemma-2b-it with varying number of seen users. The number of annotations per user remains constant except in the case with $\times 2$,” where we double the per-user annotations only for $5,000$ users, making the total number of annotations $10,000$.
Ablation study of the number of users.
We conduct an ablation study of CoPL by varying the number of users and report the performance in Fig.˜7. The performance of the model is consistent except for the case where there are only $5,000$ users in the training set. The performance with 5,000 users becomes comparable when we double the number of annotations $(2\times)$, indicating the need for a sufficient amount of annotations to capture diverse preferences.
Training reward models with GNN.
Table˜6 reports GNN accuracy on seen users and responses for test pairs excluded from the training dataset. The results demonstrate that GNN can accurately predict labels for unannotated pairs with sparse annotations. We provide the additional ablation study of message-passing in Appendix˜E.
Table˜7 examines the impact of training with GNN-based pseudo labels, allowing the model to leverage additional preference data. Although the pseudo-labeled pairs increase the dataset size, performance is slightly worse than using only user-provided annotations, suggesting that noise degrades model accuracy.
To investigate the effect of noise further, a user-specific reward model is trained on pseudo labels for a random sample of 10 users per group. The results are considerably worse than the G-Oracle, indicating that noisy labels introduce training instability. This observation aligns with Wang et al. (2024a), which notes that noisy preference labels can lead to training instability and performance degradation.
@@
Table 6: Test accuracy of the GNN. We evaluate the model using the same users from training but with annotation pairs that are not reflected in the graph.
UF-P-2 (ALL) UF-P-4 (ALL)
@@
Pseudo label $77_{\pm 0.70} 62.26$$26_{\pm 0.27}$
G-Oracle $64.53_{\pm 0.14}$ $61.52_{\pm 0.13}$
User-specific $8.09_{\pm 1.73} 55$$5.30_{\pm 3.30}$
Table 7: Accuracy of reward model trained by using a pre-trained GNN in UF-P-2/4 (ALL) with gemma-2b-it. The "pseudo-label" trains a reward model on all seen user–response pairs, with annotations provided by GNN-predicted labels. The "user-specific" refers to a BTL model trained with pseudo-labels for each user. Only 10 users per group are sampled due to computational cost.
@@ 6 Conclusion
This work demonstrates how GCF-based user embeddings enable personalization in sparse settings, but we do not extensively explore other GNN architectures that could further reduce sample complexity. Additionally, although CoPL employs a gating mechanism for user-specific expert allocation, we did not apply load-balancing loss, which induces more even activation among experts. As a result, some experts remain inactive in Fig.˜6. Future work may investigate different GNN designs and incorporate load-balancing techniques to fully leverage the potential of GNN and MoLE, respectively.
The group-wise oracle model may appear underwhelming, likely because our smaller backbone LLM struggles to capture subtle stylistic differences between responses. Larger-scale models (over 30B parameters) could better handle these nuances; however, constraints in our current setup prevent such experiments, and we defer them to future work.
Although CoPL is robust in sparse regimes compared to prior methods, it still depends on having sufficient annotation overlap to train the graph-based collaborative filtering model. In cases where the overlap is exceedingly limited, this reliance may constrain the model’s flexibility. While existing preference datasets often contain such overlap (Wang et al., 2024b; Stiennon et al., 2020; Zhang et al., 2024; Bai et al., 2022), relaxing this requirement is an important next step. A promising approach is to construct user–response graphs from semantic similarity computed with sentence embeddings or other textual similarity measures, which would extend CoPL to settings without explicit overlap.
The effectiveness of our adaptation procedure depends on the informativeness of a new user’s annotations. When annotated pairs from a user mostly involve common pairs, they contain little information about that user’s preferences, thereby degrading adaptation performance. Integrating active learning to select informative pairs for annotation could mitigate this issue and reduce sample complexity.
Finally, Fig.˜4 and Table˜A2 show that CoPL can favor majority groups under severe imbalance, even though it captures diverse preferences overall. Exploring loss reweighting is a promising direction. Methods such as focal loss (Lin et al., 2017; Subramanian et al., 2021), which increase the weight on high-error or underrepresented examples, may reduce majority bias and improve robustness.
Acknowledgements
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (RS-2024-00337955; RS-2023-00217286) and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (RS-2024-00457882, National AI Research Lab Project; RS-2019-II191906, Artificial Intelligence Graduate School Program(POSTECH)).
@@ Appendix B Method Baselines
Appendix C Related Works
Personalized alignment.
Preference learning with sparse interactions.
Appendix D Experimental Details
@@
D.1 Datasets
TL;DR.
The TL;DR dataset (Stiennon et al., 2020) contains Reddit posts alongside concise summaries and annotator IDs. Prior works (Li et al., 2022; Chen et al., 2024a) employ a modified version of this dataset by defining two simulated preference groups: one group favors shorter summaries, while the other prefers longer ones. The two groups provide different annotations for each summary pair. To focus on the most active annotators, they retain only the ten users with the highest number of annotations. We adopt the resulting set of annotation pairs from these ten users as our survey set.
Ultrafeedback-P.
@@ D.1 Datasets
UF-P-2 employs only two attributes and removes pairs that both user groups label identically, focusing on controversial cases where preferences differ. In UF-P-4, all four attributes are retained as preference dimensions, which allows for partial agreement among groups and hence increases complexity. Although Poddar et al. (2024) also excludes pairs fully agreed upon by all users, the remaining set is larger and exhibits more variety than UF-P-2.
@@
PersonalLLM.
PersonalLLM (Zollo et al., 2024) is built with 10,402 open-ended prompts that were sampled from a larger pool of 37,919 conversational questions drawn from public RLHF and preference benchmarks such as Anthropic HH-RLHF (Bai et al., 2022), NVIDIA HelpSteer (Wang et al., 2023), and RewardBench (Lambert et al., 2024). For each prompt, they used eight frontier chat models to generate a diverse response set that minimizes obvious quality gaps while covering latent preference dimensions. The resulting (prompt, response1, response2, …, response8) tuples are split into 9,402 training and 1,000 test items.
Each response is evaluated by ten strong open-source reward models with heterogeneous alignment objectives. These reward models assign scalar scores capturing distinct value dimensions for every response. Storing the full 10×8 matrix of scores per prompt provides a dense, model-agnostic preference signal that later steps can recombine to reflect arbitrary preferences. To simulate a large user base, they treat the preference of a user as a weighted ensemble over the ten reward models. The weight is sampled from a Dirichlet distribution, where varying the concentration parameter controls preference diversity.
We use $\alpha=0.1$ for Dirichlet distribution. Due to computational constraints, we simplify the dataset by selecting three responses per prompt and considering only four reward dimensions. Following Poddar et al. (2024), we remove non-controversial response pairs—those in which one response is strictly ranked below the other across all preference dimensions—to ensure the heterogeneity.
D.2 Hyper-parameters
@@
The model consists of four message-passing layers, each with user and response embeddings of dimension 512. We use Leaky ReLU as a non-linear activation function to update user and response embeddings. Training proceeds for 300 epochs using the AdamW optimizer Loshchilov (2017) with a learning rate of $1\times 10^{-4}$ and a cosine scheduler with warmup ratio $0.1$. The batch size is 1024, and all experiments are conducted on an RTX 4090 GPU.
@@
Baseline models use LoRA with rank 64. They also trained with an AdamW optimizer and a cosine scheduler with a warmup ratio $0.03$. We search the learning rate from $1\times 10^{-4},5\times 10^{-5},1\times 10^{-5},5\times 10^{-6}]$.
@@
Appendix E Additional Experimental Results
Performance under the imbalanced group distribution with UF-P-4 (AVG).
Table˜A2 reports group-wise accuracies for the four group UF-P-4 (AVG) setting under selected imbalance configurations. The results exhibit the same trend seen in the two-group setting. CoPL continues to capture diverse user preferences across all groups. As the distribution departs from the balanced 1:1:1:1 setting, the gap from the balanced baseline widens. The lower absolute accuracy of some groups is largely due to the intrinsic difficulty of their preferences rather than the imbalance itself. This interpretation is supported by the G-Oracle. Fig.˜A6 visualizes the learned user embeddings. The embeddings form well-separated clusters aligned with group identities even under strong imbalance, which suggests that the representation remains stable, although predictive performance on minority groups drops.
Ablation study of the number of users.
Fig.˜A3 shows that CoPL performs robustly across different expert counts. This indicates that a moderate number of experts is generally sufficient to capture diverse user preferences.
Performance with large-scale LLM.
To assess scalability, we instantiate CoPL with gemma-2-27B-it (Team et al., 2024b) and evaluate on UF-P-4 (ALL) and PersonalLLM (ALL). We use a single seed due to hardware limits and compare with VPL, the strongest baseline. As shown in Table˜A3, CoPL surpasses VPL on both datasets, indicating that the gains carry over to larger model scales. These results support the scalability of CoPL beyond the settings used in the main experiments.
Ablation study of message-passing.
Inspired by the previous work He et al. (2020) in recommendation systems, we first omit the non-linear activation and feature transformation matrix used in Eq.˜2, and also investigate the effectiveness of negative edges. As shown in  Table˜A4, incorporating negative edges consistently improves accuracy. Notably, our proposed message-passing achieves the highest accuracy, highlighting both the effectiveness of our message-passing operation and the advantage of modeling negative edges.
TL;DR UF-P-2 UF-P-4 PersonalLLM
ALL AVG ALL AVG ALL AVG ALL AVG
Seen
G-Oracle $7.21_{\pm 0.28} 77$$7.21_{\pm 0.28}$ $6.80_{\pm 0.17} 66$$6.80_{\pm 0.17}$ $62.17_{\pm 0.09}$ $62.17_{\pm 0.09}$ N/A N/A
Uniform $49.39_{\pm 0.52}$ $49.39_{\pm 0.52}$ $61.96_{\pm 0.07}$ $61.96_{\pm 0.07}$ $56.80_{\pm 0.12}$ $56.80_{\pm 0.12}$ $63.64_{\pm 0.30}$ $63.64_{\pm 0.30}$
I2E $49.40_{\pm 0.77}$ $49.66_{\pm 0.31}$ $62.10_{\pm 0.28}$ $61.43_{\pm 0.23}$ $57.90_{\pm 0.21}$ $58.50_{\pm 0.09}$ $66.40_{\pm 0.38}$ $65.86_{\pm 0.12}$
[…]
VPL $49.14_{\pm 0.72}$ $49.17_{\pm 0.67}$ $62.39_{\pm 0.10}$ $62.59_{\pm 0.24}$ $58.87_{\pm 0.25}$ $57.55_{\pm 1.00}$ $70.55_{\pm 0.16}$ $66.18_{\pm 0.01}$
PAL $49.57_{\pm 0.09}$ $49.75_{\pm 0.27}$ $62.59_{\pm 0.06}$ $62.47_{\pm 0.13}$ $57.17_{\pm 0.22}$ $56.27_{\pm 0.13}$ $66.46_{\pm 0.49}$ $65.43_{\pm 0.43}$
CoPL $\mathbf{97.85}_{\pm 0.07}$ $\mathbf{97.88}_{\pm 0.01}$ $\mathbf{63.90}_{\pm 0.07}$ $\mathbf{63.48}_{\pm 0.13}$ $\mathbf{62.90}_{\pm 0.05}$ $\mathbf{61.93}_{\pm 0.02}$ $\mathbf{74.87}_{\pm 0.19}$ $\mathbf{74.76}_{\pm 0.01}$
Unseen
G-Oracle $7.54_{\pm 0.49} 77$$7.54_{\pm 0.49}$ $67.43_{\pm 0.65}$ $67.43_{\pm 0.65}$ $62.01_{\pm 0.04}$ $62.01_{\pm 0.04}$ N/A N/A
Uniform $49.03_{\pm 0.76}$ $49.03_{\pm 0.76}$ $62.23_{\pm 0.06}$ $62.23_{\pm 0.06}$ $_{\pm 0.27} 57.02±0.275$$_{\pm 0.27}$ $63.30_{\pm 0.08}$ $63.30_{\pm 0.08}$
I2E $49.64_{\pm 0.98}$ $49.56_{\pm 0.49}$ $62.62_{\pm 0.95}$ $61.88_{\pm 0.21}$ $57.62_{\pm 0.92}$ $58.12_{\pm 0.98}$ $65.75_{\pm 0.38}$ $65.74_{\pm 0.37}$
[…]
VPL $49.07_{\pm 0.65}$ $48.92_{\pm 0.72}$ $62.69_{\pm 0.99}$ $63.67_{\pm 0.12}$ $58.49_{\pm 1.22}$ $56.85_{\pm 0.84}$ $69.93_{\pm 0.33}$ $65.72_{\pm 0.42}$
PAL $49.71_{\pm 0.44}$ $49.68_{\pm 0.34}$ $63.08_{\pm 0.73}$ $62.52_{\pm 0.58}$ $57.15_{\pm 0.48}$ $56.44_{\pm 0.67}$ $66.57_{\pm 0.08}$ $65.92_{\pm 0.25}$
CoPL $\mathbf{97.95}_{\pm 0.15}$ $\mathbf{98.19}_{\pm 0.06}$ $\mathbf{64.08}_{\pm 0.71}$ $\mathbf{64.38}_{\pm 1.00}$ $\mathbf{62.77}_{\pm 1.32}$ $\mathbf{62.08}_{\pm 0.64}$ $\mathbf{74.84}_{\pm 0.18}$ $\mathbf{75.64}_{\pm 0.05}$
Table A1: Accuracy of reward models on unseen annotated pairs. The results report performance on Seen users encountered during training and on Unseen users, which consist of 100 new users evenly distributed across preference groups. Unseen users provide 8 annotations under TL;DR/UF-P-2 (ALL/AVG) and 16 annotations under UF-P-4/PersonalLLM (ALL/AVG). Bold represents the best result, except for G-Oracle. N/A indicates that training reward models for each group is infeasible for PersonalLLM, as this dataset does not clearly partition users into discrete groups. All experiments run on three seeds. These results are based on gemma-7b-it.
(a) I2E
[…]
(c) VPL
(d) PAL
(e) CoPL
Figure A1: T-SNE visualization of seen user embeddings in UF-P-4 (AVG) with gemma-2b-it. Points are colored by their preference group. Our method clusters users in the same group more effectively, whereas other baselines fail to cluster users by their preference groups in the user embedding space.